a number of scientists between the 10th and the 17th centuries (sahl, snellius, descartes) concluded that the following formula should relate the angle of incidence with the angle of refraction. Sini = sinr/n. Explain ?


1
Expert's answer
2021-04-23T07:27:27-0400

"sini = \frac{sinr}{n}"

It explains that whenever the light goes from one medium to another medium its speed changes due to the characteristics of that medium. Such property of medium is measured by a quantity called reflective index n.

Suppose if the light goes from air to any other denser medium of refractive index n. Then considering the refractive index of air n=1, we found

n (refractive index of denser medium) "= \frac{sini}{sinr}"

If the speed of light in air is c and in the denser medium of refractive index is v, then the above law can be written as

"\frac{c}{v}=\frac{sini}{sinr}"

c>v hence i>r